Classify each of the following reactions?

1. Mg + HCL =MgCL2+H2

2.P4O10+C = P4 + CO2

3.AgNO3 +NaBr = AgBr + NaNO3

4.Ba(OH)2+H2SO4 =BaSO4+H2O

5.CL2+ KBr = Br2+KCL



1
Expert's answer
2021-04-14T07:06:09-0400

Mg + HCL =MgCL2+H2

oxidation-reduction exchange reaction

P4O10+C = P+ CO2

Single displacement reaction

AgNO+NaBr = AgBr + NaNO3

Precipitation reaction

Ba(OH)2+H2SO=BaSO4+H2O

Acid-Base reaction

CL2+ KBr = Br2+KCL

Single displacement reaction